# Toward a Cognitive-Affective-Systemic Framework for Art and Sustainability

### v2

@@ Abstract
Complexity; art and sustainability; logomotion; cognitive–affective–systemic framework; art and science
@@ 1 Introduction
Scientific communication through data, graphs, and models makes these dynamics visible but seldom affective. Understanding alone does not ensure transformation. To confront planetary crises, knowledge must become embodied, affective, and participatory. Art occupies this crucial space. Through sensory experience, imagination, and affective resonance, artistic practice can cultivate ecological awareness that transcends data, activating empathy, curiosity, and care. As Duxbury, Kangas, and de Coninck (2017) note, the arts play a vital role in sustainable development by “raising awareness and catalysing action about sustainability and climate change.” Artistic practice thus becomes not only representational but epistemic: a way of knowing through making and feeling.
@@ 4.1 SPill: The avalanche of antimicrobial resistance
In order to raise the awareness about the AMR issue, we employed the physical mechanism behind the avalanche effect—the sandpile model—to construct the kinetic sound installation SPill. The sandpile model was proposed in 1987 by physicists Bak, Tang, and Wiesenfeld (1987) to illustrate the concept of self-organized criticality—how complex, scale-free behavior can emerge spontaneously in nature. In the model, grains of sand are dropped one by one onto a grid. When a site exceeds a critical threshold, it topples, distributing grains to its neighbors. This can trigger a cascade of further toppling. Over time, the system naturally evolves toward a critical state where avalanches of all sizes occur, following a power-law distribution without external fine-tuning.
@@
Many viewers described being captivated by the simultaneous fragility and inevitability of the cascading motion. Others were drawn to the visual aspect of the artwork—the vibrant colours, the avalanche, and the incredibly realistic avalanche sound generated by the movement of pills. Such curiosity and desire to participate motivated audiences to contribute their unwanted medicines to the installation. All the pills used in the work were ultimately recycled through pharmacies following proper disposal procedures.
@@ 4.2 Echoes of the Land: The embodied perception of anthropogenic seismicity
The interactive installation Echoes of the Land transforms this phenomenon into a multisensory environment that merges scientific modelling with participatory experience. Its design is grounded in the spring-block model of earthquake recurrence (Burridge and Knopoff 1967; Brown, Scholz, and Rundle 1991), which simulates tectonic stick-slip dynamics through a grid of masses connected by springs on a frictional surface. As stress accumulates gradually, it is released in sudden cascades once a critical threshold is surpassed—an archetype of self-organized criticality in geophysical systems.
Figure 3: (a) The spring-block model of earthquake recurrence proposed by Brown, Scholz, and Rundle (1991). (b) Adapted $5\times 5$ spring-block system used in the installation Echoes of the Land, enclosed within a movable frame for human interaction.
@@
In the artwork, this model is re-imagined as a $5\times 5$ spring-block ensemble enclosed within a movable square frame. Each block is connected by springs that allow limited horizontal motion, while the frame rests on caster wheels, permitting smooth, multidirectional movement. Participants can push, pull, or rotate the frame, generating stress within the network. When frictional limits are exceeded, the blocks slip, releasing motion that propagates through the array like seismic waves. Embedded accelerometers in each block detects these micro-movements, triggering flashes of white light through semi-translucent 3D-printed covers derived from digital terrain data Figure 5. The visual effect suggests fragments of shifting landscape—an Earth responding to the pressure of human touch.
@@
As participants interact, they effectively generate their own earthquake data, enacting the feedback loop between human agency and geological response. The installation’s tactile, visual, and sonic feedback transforms spectators into co-agents, allowing them to feel the thresholds and tensions that define the Earth system. Many participants reported sensing an uncanny awareness of responsibility as their small gestures triggered large-scale visual and sonic cascades, an affective embodiment of complexity itself. This uncanniness was intentionally created as a “reflectaphor” within the narrative structure. As described by Jennings (1996), it generates “an irresolvable tension between similarity and differences.” We utilized the reflectaphor to pique the audience’s interest, curiosity, and emotions, with the intention of directing their attention towards the story conveyed by the works. Further detailed technical description and aesthetic analysis are published elsewhere (Liu, Hao, and Xie 2025).
Echoes of the Land has been exhibited internationally at venues including Ars Electronica (2025), ISEA (2025), Bauhaus Museum (Weimar Kunstfest 2025), LEV Festival (2025), and IRCAM (2024), where it received enthusiastic critical reception. During these exhibitions, audiences not only experienced the multisensory impact of the installation but also learned about the scientific principles behind earthquakes through on-site demonstrations and explanatory materials. The work thus functioned as a form of science communication, translating seismological modelling into embodied and intuitive experience. Many visitors remarked on the ingenuity of the interactive design, noting how the piece seamlessly integrated physical mechanics, data translation, and immersive media to reveal the hidden dynamics of the Earth system.
Across these presentations, the work prompted spontaneous dialogue among audiences, who connected the piece to analogous local phenomena in their own regions, underscoring the global resonance of induced seismicity.
@@ Acknowledgements
No potential conflict of interest was reported by the author(s).

### v3

# Art Practice for Sustainability: A Cognitive-Affective-Systemic Framework

@@ Abstract
Complexity; art and sustainability; visualization; media art; art and science
@@ 1 Introduction
Scientific communication through data, graphs, and models makes these dynamics visible but seldom affective. Understanding alone does not ensure transformation. To confront planetary crises, knowledge must become embodied, affective, and participatory. Art occupies this crucial space. Through sensory experience, imagination, and affective resonance, artistic practice can cultivate ecological awareness that transcends data, activating empathy, curiosity, and care. As Duxbury et al. (2017) note, the arts play a vital role in sustainable development by “raising awareness and catalysing action about sustainability and climate change.” Artistic practice thus becomes not only representational but epistemic: a way of knowing through making and feeling.
@@ 3.1 The three dimensions
Alt Text: A conceptual diagram showing the relationship between art, awareness, and engagement/action. Art sits at the top as an interface bridging cognition, emotion, and systemic understanding. Three interlinked arrows form a cyclic loop between cognitive awareness, emotional awareness, and systemic awareness, while a vertical flow connects art practice to collective reflection and action.
@@ 4.1 SPill: The avalanche of antimicrobial resistance
Alt Text: (a) A kinetic installation featuring a transparent container filled with colorful recycled pills forming a slope. A mechanical lifting system continuously moves the pills upward, creating periodic avalanche-like cascades. (b) Close-up view of unpackaged recycled pills of varied colors and sizes arranged for use in the installation, highlighting pharmaceutical waste as artistic material.
In order to raise the awareness about the AMR issue, we employed the physical mechanism behind the avalanche effect—the sandpile model—to construct the kinetic sound installation SPill. The sandpile model was proposed in 1987 by physicists Bak et al. (1987) to illustrate the concept of self-organized criticality—how complex, scale-free behavior can emerge spontaneously in nature. In the model, grains of sand are dropped one by one onto a grid. When a site exceeds a critical threshold, it topples, distributing grains to its neighbors. This can trigger a cascade of further toppling. Over time, the system naturally evolves toward a critical state where avalanches of all sizes occur, following a power-law distribution without external fine-tuning.
@@
The cascading motion captivates the viewers with a sense of simultaneous fragility and inevitability of the cascading motion. The visual aspect of the artwork also draws their attention—the vibrant colours, the avalanche, and the incredibly realistic avalanche sound generated by the movement of pills. Such curiosity and desire to participate motivated audiences to contribute their unwanted medicines to the installation. All the pills used in the work were ultimately recycled through pharmacies following proper disposal procedures.
@@ 4.2 Echoes of the Land: The embodied perception of anthropogenic seismicity
The interactive installation Echoes of the Land transforms this phenomenon into a multisensory environment that merges scientific modelling with participatory experience (Figure 4). Its design is grounded in the spring-block model of earthquake recurrence (Burridge and Knopoff 1967; Brown et al. 1991), which simulates tectonic stick-slip dynamics through a grid of masses connected by springs on a frictional surface. As stress accumulates gradually, it is released in sudden cascades once a critical threshold is surpassed—an archetype of self-organized criticality in geophysical systems.
Figure 3: (a) The spring-block model of earthquake recurrence proposed by Brown et al. (1991). (b) Adapted $5\times 5$ spring-block system used in the installation Echoes of the Land, enclosed within a movable frame for human interaction.
Alt Text: (a) A schematic of the original spring-block model showing an array of interconnected blocks and springs simulating tectonic stress and slip. (b) A physical $5\times 5$ grid of small square blocks joined by springs within a movable square frame, representing a tangible adaptation of the earthquake model for audience interaction
@@
Alt Text: A large interactive installation combining a grid of spring-blocks on a movable frame with immersive projection and sound. Participants push or move the frame, triggering lights and seismic-like sounds. Projected images of human infrastructure and environmental destruction appear on surrounding walls, linking human action to seismic consequences.
@@
Alt Text: (a) Detailed view of the spring-block structure resembling an undulating terrain, emphasizing the sculptural surface and interlinked motion system. (b) Each block contains a small accelerometer; when movement occurs, bright white lights flash through semi-translucent covers, visualizing the propagation of simulated seismic waves..
In the artwork, this model is re-imagined as a $5\times 5$ spring-block ensemble enclosed within a movable square frame. Each block is connected by springs that allow limited horizontal motion, while the frame rests on caster wheels, permitting smooth, multidirectional movement. Audience can push, pull, or rotate the frame, generating stress within the network. When frictional limits are exceeded, the blocks slip, releasing motion that propagates through the array like seismic waves. Embedded accelerometers in each block detects these micro-movements, triggering flashes of white light through semi-translucent 3D-printed covers derived from digital terrain data Figure 5. The visual effect suggests fragments of shifting landscape—an Earth responding to the pressure of human touch.
@@
As audience interact, they effectively generate their own earthquake data, enacting the feedback loop between human agency and geological response. The installation’s tactile, visual, and sonic feedback transforms spectators into co-agents, allowing them to feel the thresholds and tensions that define the Earth system. This design intents to create an uncanny awareness of responsibility as the audience’s small gestures triggered large-scale visual and sonic cascades, an affective embodiment of complexity itself. This uncanniness was intentionally created as a “reflectaphor” within the narrative structure. As described by Jennings (1996), it generates “an irresolvable tension between similarity and differences.” We utilized the reflectaphor to pique the audience’s interest, curiosity, and emotions, with the intention of directing their attention towards the story conveyed by the works. Further detailed technical description and aesthetic analysis are published elsewhere (Authors 2025).
Echoes of the Land has been exhibited internationally at venues including Ars Electronica (2025), ISEA (2025), Bauhaus Museum (Weimar Kunstfest 2025), LEV Festival (2025), and IRCAM (2024), where it received enthusiastic critical reception. During these exhibitions, audiences not only experienced the multisensory impact of the installation but also learned about the scientific principles behind earthquakes through on-site demonstrations and explanatory materials. The work thus functioned as a form of science communication, translating seismological modelling into embodied and intuitive experience.
Across these presentations, the work was designed to prompt spontaneous dialogue among audiences, connecting the piece to analogous local phenomena in their own regions, underscoring the global resonance of induced seismicity.
@@ Acknowledgements
The author reports no potential conflict of interest.